The length of a rectangle is 3 yd less than twice the width, and the area of the rectangle is 65 yd2 . Find the dimensions of the rectangle.
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = 2x - 3
Area = Width x Length
65 = x(2x - 3)
65 = 2x^2 - 3x
Sort into ax^2 + bx + c = 0 form.
2x^2 - 3x - 65 = 0
Factorise:
(2x - 13)(x + 5) = 0
x + 5 = 0
x + -5 (Disregard as -ve)
2x - 13 = 0
2x = 13
x = 6.5
Width = 6.5 yds
Length = 10 yds.
Hope this helps :-)